# Solver-in-the-Loop: MDP-Based Benchmarks for Self-Correction and Behavioral Rationality in Operations Research

### v1

@@ 2.1 OR-Debug-Bench: Data Organization & Metrics
Difficulty Levels Easy / Medium / Hard / Expert
@@ 3.1 Saboteur-based Problem Generation
Type Name IIS Size Difficulty
A Direction Flip 2–3 Easy
B RHS Miscalculation 3–5 Medium
C Upper Bound Conflict 2–3 Easy
D Lower Bound Conflict 2–4 Easy
E Resource Over-allocation 5–8 Hard
F Capacity Violation 5–7 Hard
G Flow Imbalance 6–10 Hard
H Multi-constraint Conflict 8–12 Expert
I Composite Error 10–15 Expert
@@
Table 7: Difficulty calibration for OR-Debug-Bench. Each level targets specific IIS sizes and expected SFT baseline performance.
Level IIS Range Error Types Target SFT RR@5
Easy 2–4 A, B, C, D $\geq$90%
Hard 5–10 E, F, G 70–85%
Expert 8–15 H, I 45–70%
@@ 5.1 Experimental Setup
5.2 Main Results
Table 11: OR-Debug-Bench results on representative models (450 stratified samples). Full results for all 26 models in Appendix E.
@@ 5.2 Main Results
Qwen3-8B-GRPO 100% 95.3% 62.4% 2.25
Qwen3-8B-Curriculum 100% 94.0% 61.7% 2.22
Qwen3-8B-DAPO 100% 93.8% 60.4% 2.31
Qwen3-8B-SFT 99.8% 93.1% 60.8% 2.34
@@
Per-Error-Type Performance. Domain-specific training provides larger gains on harder problems: +13.3% on Types E–G (94.7% vs 81.4%) and +20.8% on Types H–I (89.1% vs 68.3%). All models achieve >>90% on Types A–D, confirming the baseline difficulty calibration. Full breakdown appears in Appendix E.2.
@@ 6 Related Work
OR Benchmarks. Table 2 compares OR-Debug-Bench and OR-Bias-Bench with existing OR benchmarks. NL4Opt (Ramamonjison et al., 2023) introduced the task of translating natural language to linear program formulations. Subsequent work explored both prompt-based approaches using in-context learning (Xiao et al., 2024; AhmadiTeshnizi et al., 2024; Bertsimas and Margaritis, 2024) and learning-based methods fine-tuning on domain data (Yang et al., 2024; Huang et al., 2025). Other recent efforts include ORQA (Mostajabdaveh et al., 2025) for OR reasoning evaluation, LLMOPT (Jiang et al., 2024), and agent-based systems (Zhang and Luo, 2025; Thind et al., 2025). OptMATH (Lu et al., 2025) proposed bidirectional data synthesis for scalable training data generation. MAMO (Huang et al., 2024) further contributed benchmarks spanning easy and complex LP instances with solver integration. PILOT-Bench (Ao et al., 2026b) evaluates LLM workflow execution under probabilistic tool failures and varying instruction quality. However, all existing benchmarks evaluate static, one-shot formulation accuracy: a model produces code, and evaluation checks correctness without iterative refinement or solver feedback. Recent work has extended formulation benchmarks to dynamic programming: DP-Bench (Zhou et al., 2025) introduces 132 textbook-level DP problems and shows that even SOTA models struggle with stochastic transitions, achieving only 59.1% accuracy. Our work addresses a complementary gap: evaluating how models respond to solver feedback and iteratively correct their formulations, focusing on LP infeasibility debugging rather than one-shot formulation.
@@
Behavioral Rationality in LLMs. Recent work has examined LLMs as simulators of human decision-making in OM (Zhang et al., 2025b). AIM-Bench (Zhao et al., 2025) documented the “pull-to-center” phenomenon in LLM inventory decisions, where models regress toward mean predictions regardless of the optimal decision. This finding echoes classical behavioral OR on human decision-making under uncertainty (Schweitzer and Cachon, 2000). Our OR-Bias-Bench benchmark extends AIM-Bench with explicit ID/OOD splits and demonstrates that curriculum learning can mitigate these biases, achieving 48% bias reduction on OOD scenarios.
@@ Appendix A Execution Examples
A.1 Success Case: Type C Error (Upper Bound Conflict)
@@
A.2 Challenge Case: Composite Error (Type I)
@@ A.4 Additional Success Examples
Example: Type E (Resource Over-allocation). A transportation problem where demand exceeds supply:
@@ B.6 Evaluation Protocol
Stratified Sampling. To ensure balanced evaluation across difficulty levels:
  • 180 samples from Types A–D (Easy)
  • 158 samples from Types E–G (Hard)
  • 112 samples from Types H–I (Expert)
@@ C.1 Saboteur Generation
Type E: Resource Over-allocation. This error increases demand constraints beyond available capacity:
Algorithm 4 Sabotage: Resource Over-allocation (Type E)
@@ C.6 Error Type Examples
Type B: RHS Miscalculation.
@@
m.addConstr(demand <= 100, "demand_limit")
# Demand cannot exceed 100
@@
m.addConstr(demand <= 50, "demand_limit")
# Insufficient to meet requirements
The reduced RHS makes it impossible to satisfy downstream demand constraints.
Type C: Upper Bound Conflict.
@@
m.addConstr(x[0] <= 40, "capacity")
# When minimum requirement is x[0] >= 60
The tightened upper bound conflicts with a lower bound on the same variable.
Type D: Lower Bound Conflict.
@@
m.addConstr(x[0] >= 100, "non_negative")
# When maximum capacity is x[0] <= 80
The inflated lower bound exceeds the variable’s upper limit.
Type E: Resource Over-allocation.
# Original (feasible)
m.addConstr(quicksum(demand[i] for i in I)
            <= 90, "total_demand")
@@
m.addConstr(quicksum(demand[i] for i in I)
            <= 120, "total_demand")
# When total capacity is only 100
Individual demand constraints sum to more than available capacity.
Type F: Capacity Violation.
# Original (feasible)
m.addConstr(production <= 100, "capacity")
@@
m.addConstr(production <= 60, "capacity")
# When minimum production requirement is 80
The capacity constraint is too restrictive for minimum requirements.
Type G: Flow Imbalance.
# Original (feasible)
m.addConstr(inflow[j] - outflow[j] == 0,
[…]
@@
m.addConstr(inflow[j] - outflow[j] == 50,
[…]
# Creates impossible flow requirement
The modified balance constraint cannot be satisfied by any feasible flow.
Type H: Multi-constraint Conflict. Multiple constraints interact to create infeasibility:
# Sabotaged constraints
m.addConstr(x + y >= 80, "c1")
m.addConstr(x <= 30, "c2")
m.addConstr(y <= 40, "c3")
# No solution: max(x+y) = 70 < 80
Resolving the conflict requires understanding how all three constraints interact.
Type I: Composite Error. Combines multiple error types—e.g., a direction flip (Type A) that creates a cascade of bound conflicts (Type C/D):
# Multiple sabotaged constraints
m.addConstr(x + y <= 50, "flipped")    # Type A
m.addConstr(x >= 30, "bound1")          # Type D
m.addConstr(y >= 30, "bound2")          # Type D
# Infeasible: x + y >= 60 > 50
@@ C.11 Difficulty Calibration Procedure
We calibrate difficulty levels through iterative testing against a baseline SFT model. The procedure ensures each difficulty level provides meaningful differentiation.
@@
    Initial parameter selection: Set IIS size ranges based on error type complexity.
@@
    Baseline evaluation: Run SFT model on 100 problems per configuration.
@@
    Target verification: Check if RR@5 falls within target range.
@@
    Parameter adjustment: If too easy (RR@5 >> target), increase IIS size or use harder error types. If too hard (RR@5 << target), decrease IIS size or use easier error types.
@@
    Iteration: Repeat until convergence (typically 2–3 iterations).
@@
Table 23: Difficulty calibration results after iterative refinement.
Level IIS Range Target RR@5 Observed RR@5 Iterations
Easy 2–4 $\geq$90% 93.1% 1
Hard 5–10 70–85% 78.4% 2
Expert 8–15 45–70% 62.2% 3
@@
  • >>90% (Easy): Ensures baseline models can solve simple problems, establishing floor performance.
  • 70–85% (Hard): Provides room for improvement while remaining tractable.
  • 45–70% (Expert): Challenges advanced models without being impossibly difficult.
@@
The Expert level required 3 calibration iterations because initial IIS sizes of 10–20 resulted in only 35% RR@5, below our utility threshold. Reducing to 8–15 achieved the target range while maintaining genuine difficulty.
@@ H.3 Scaling with Problem Difficulty
Easy B (87%), C (76%) 81.4% Clear diagnostic signals
Medium H (65%), I (64%), E (61%), D (57%) 61.6% Multiple constraint interactions
Hard A (54%), F (46%), G (45%) 48.1% Semantic reasoning required
@@
  • Type B is easiest, not Type A: Bound violations (Type B) produce clear diagnostic signals with 87% average success. Type A (objective direction), despite being a single-constraint error, averages only 54%—models struggle with objective function semantics.
  • Flow and capacity are hardest: Types F (flow balance, 46%) and G (capacity violations, 45%) require reasoning about multiple interacting constraints, explaining the difficulty gap.
@@ H.4 Recommendations for Practitioners
    Apply difficulty-adaptive resources: Allocate more compute to Types A, F, G (Hard, 48% avg RR@5) than Types B, C (Easy, 81% avg).

### v2

@@ 2.1 OR-Debug-Bench: Data Organization & Metrics
Difficulty Levels Easy / Medium / Hard
@@ 3.1 Saboteur-based Problem Generation
Type Name Difficulty
A Direction Flip Hard
B Variable Type Error Easy
C Coefficient Modification Easy
D Contradicting Constraint Hard
E Multi-Constraint Conflict Hard
F Hidden Dependency Hard
G Cascading Conflict Hard
H IIS-Incomplete Medium
I Optimal Selection Medium
@@
Table 7: Difficulty calibration for OR-Debug-Bench. Levels are defined by baseline API model performance.
Level Error Types Baseline RR@5
Easy B, C $\geq$85%
Medium H, I 70–85%
Hard A, D, E, F, G <<70%
@@ 5.1 Experimental Setup
5.2 Trained Local Models
[…]
@@ 5.2 Trained Local Models
Qwen3-8B
     GRPO 100% 95.3% 62.4% 2.25
     Curriculum 100% 94.0% 61.7% 2.22
     DAPO 100% 93.8% 60.4% 2.31
     SFT 99.8% 93.1% 60.8% 2.34
Llama-3.1-8B
     GRPO 100% 97.3% 60.3% 1.82
     SFT 100% 97.1% 60.4% 1.85
DeepSeek-R1-Distill-Qwen3-8B
     GRPO 99.6% 88.9% 52.2% 3.30
     SFT 99.3% 88.7% 52.2% 3.30
Table 12: API model baselines on OR-Debug-Bench (no training of LLM models). Full results for all 22 API models in Appendix E.
Model $R$ RR@5 DA Steps
@@
Table 11 presents results for trained local models. All three model families (Qwen, Llama, DeepSeek) benefit from the same training pipeline: supervised fine-tuning on expert trajectories followed by GRPO reinforcement learning with solver-verified rewards. Llama-3.1-8B-GRPO achieves the highest RR@5 (97.3%) with fastest repair (1.82 steps). Qwen3-8B-GRPO achieves the highest DA (62.4%), indicating more accurate root-cause identification. DeepSeek-R1-Distill-Qwen3-8B, starting from a reasoning-distilled base, shows competitive performance (88.9% RR@5).
Training Effectiveness: Across all model families, GRPO training consistently improves over SFT: +0.2–2.2% RR@5 improvement. The consistent gains across diverse architectures demonstrate the generalizability of our training approach.
5.3 API Model Baselines
Table 12 presents API model baselines evaluated without any training. The best API models (o4-mini, claude-sonnet-4) achieve 86.2% RR@5, substantially below the trained 8B models.
@@ 5.3 API Model Baselines
Per-Error-Type Performance. Domain-specific training provides larger gains on harder problems (A, D–G): +9.6% average (94.4% vs 84.8%). Easy types (B, C) show smaller gains (+3.0%) as baselines already exceed 95%. Medium types (H, I) improve by +14.0% (95.0% vs 81.0%). Full breakdown appears in Appendix E.2.
@@ 6 Related Work
OR Benchmarks. Table 2 compares OR-Debug-Bench and OR-Bias-Bench with existing OR benchmarks. NL4Opt (Ramamonjison et al., 2023) introduced the task of translating natural language to linear program formulations. Subsequent work explored both prompt-based approaches using in-context learning (Xiao et al., 2024; AhmadiTeshnizi et al., 2024; Bertsimas and Margaritis, 2024) and learning-based methods fine-tuning on domain data (Yang et al., 2024; Huang et al., 2025). Other recent efforts include ORQA (Mostajabdaveh et al., 2025) for OR reasoning evaluation, LLMOPT (Jiang et al., 2024), and agent-based systems (Zhang and Luo, 2025; Thind et al., 2025). OptMATH (Lu et al., 2025) proposed bidirectional data synthesis for scalable training data generation. MAMO (Huang et al., 2024) further contributed benchmarks spanning easy and complex LP instances with solver integration. PILOT-Bench (Ao et al., 2026c) evaluates LLM workflow execution under probabilistic tool failures and varying instruction quality. However, all existing benchmarks evaluate static, one-shot formulation accuracy: a model produces code, and evaluation checks correctness without iterative refinement or solver feedback. Recent work has extended formulation benchmarks to dynamic programming: DP-Bench (Zhou et al., 2025) introduces 132 textbook-level DP problems and shows that even SOTA models struggle with stochastic transitions, achieving only 59.1% accuracy. Our work addresses a complementary gap: evaluating how models respond to solver feedback and iteratively correct their formulations, focusing on LP infeasibility debugging rather than one-shot formulation.
@@
Behavioral Rationality in LLMs. Recent work has examined LLMs as decision-makers in OM contexts, including demand simulation (Zhang et al., 2025b) and sequential evaluation under limited human feedback (Ao et al., 2026a). AIM-Bench (Zhao et al., 2025) documented the “pull-to-center” phenomenon in LLM inventory decisions, where models regress toward mean predictions regardless of the optimal decision. This finding echoes classical behavioral OR on human decision-making under uncertainty (Schweitzer and Cachon, 2000). Our OR-Bias-Bench benchmark extends AIM-Bench with explicit ID/OOD splits and demonstrates that curriculum learning can mitigate these biases, achieving 48% bias reduction on OOD scenarios.
@@ Appendix A Execution Examples
A.1 Success Case: Type C Error (Coefficient Modification)
@@
A.2 Challenge Case: Optimal Selection (Type I)
@@ A.4 Additional Success Examples
Example: Type E (Multi-Constraint Conflict). A transportation problem where demand exceeds supply:
@@ B.6 Evaluation Protocol
Stratified Sampling. To ensure balanced evaluation across difficulty levels (50 samples per error type):
  • 100 samples from Types B, C (Easy)
  • 100 samples from Types H, I (Medium)
  • 250 samples from Types A, D, E, F, G (Hard)
@@ C.1 Saboteur Generation
Type E: Multi-Constraint Conflict. This error creates interlocked constraints that both must be fixed:
Algorithm 4 Sabotage: Multi-Constraint Conflict (Type E)
@@ C.6 Error Type Examples
Type B: Variable Type Error.
@@
x = m.addVar(vtype=GRB.INTEGER, ub=10, name="x")
@@
x = m.addVar(vtype=GRB.BINARY, name="x")
m.addConstr(x >= 2, "forcing")
# Binary variable cannot be >= 2
The variable type change combined with a forcing constraint creates infeasibility.
Type C: Coefficient Modification.
@@
m.addConstr(20*x + 30*y <= 100, "capacity")
# Scaled coefficients make constraint unsatisfiable
The modified coefficients make the constraint impossible to satisfy with existing bounds.
Type D: Contradicting Constraint.
@@
m.addConstr(x + y >= 150, "conflicting")
# Directly contradicts upper bound
The added constraint directly contradicts existing constraints.
Type E: Multi-Constraint Conflict.
@@
m.addConstr(x + y <= 50, "e1")
m.addConstr(x + y >= 100, "e2")
# Both constraints cannot be satisfied
Interlocked constraints require fixing multiple constraints to restore feasibility.
Type F: Hidden Dependency.
@@
aux = m.addVar(name="aux")
m.addConstr(aux == x + y, "def_aux")
m.addConstr(aux >= 200, "root_cause")  # Hidden
m.addConstr(x + y <= 100, "symptom")   # Shows in IIS
The root cause (aux $\geq$ 200) is not directly visible in the IIS.
Type G: Cascading Conflict.
@@
m.addConstr(x <= 30, "g1")   # Initial IIS
m.addConstr(x >= 50, "g2")   # Hidden until g1 fixed
m.addConstr(x <= 100, "g3")  # Original bound
Fixing the first conflict reveals another; requires understanding the cascade.
Type H: IIS-Incomplete.
# Sabotaged (infeasible)
x.LB = 80                    # Root cause (bound)
m.addConstr(x + y <= 50, "symptom")
# IIS shows symptom constraint, not the bound
The IIS shows the symptom constraint, but the root cause is a variable bound.
Type I: Optimal Selection.
# Sabotaged (infeasible)
m.addConstr(x >= 60, "lower")
m.addConstr(x <= 40, "upper")
# Multiple fixes possible, different OP impacts
Multiple repairs restore feasibility, but only one preserves the original optimal objective.
@@ C.11 Difficulty Calibration Procedure
We calibrate difficulty levels through iterative testing against baseline API models. The procedure ensures each difficulty level provides meaningful differentiation.
@@
    Initial grouping: Group error types by semantic complexity and expected reasoning requirements.
@@
    Baseline evaluation: Run API models on 50 problems per error type.
@@
    Target verification: Check if average RR@5 falls within target range for each group.
@@
    Group adjustment: Reassign error types between difficulty levels based on observed performance.
@@
    Iteration: Repeat until stable groupings (typically 2–3 iterations).
@@
Table 24: Difficulty calibration based on baseline API model performance.
Level Error Types Target RR@5 Observed RR@5
Easy B, C $\geq$85% 90.5%
Medium H, I 70–85% 78.5%
Hard A, D, E, F, G <<70% 59.0%
@@
  • $\geq$85% (Easy): Ensures baseline models can solve simple problems, establishing floor performance. Types B and C fall in this category.
  • 70–85% (Medium): Moderate difficulty with room for improvement. Types H and I fall here.
  • <<70% (Hard): Challenges models significantly while remaining tractable. Types A, D, E, F, and G require multi-step reasoning.
@@
Difficulty calibration was based on empirical baseline API model performance rather than theoretical metrics, as we found accuracy correlates more strongly with error type semantics than with other factors.
@@ H.3 Scaling with Problem Difficulty
Easy B (95%), C (86%) 90.5% Clear diagnostic signals
Medium H (82%), I (75%) 78.5% Multi-step constraint interactions
Hard A (61%), D (63%), E (69%), F (49%), G (53%) 59.0% Semantic reasoning required
@@
  • Type B is easiest: RHS miscalculation (Type B) produces clear diagnostic signals with 95% average success. Type A (direction flip), despite appearing simple, averages only 61%—flipping constraint directions creates conflicts with multiple existing constraints.
  • Types F and G are hardest: Types F (capacity violation, 49%) and G (flow imbalance, 53%) require reasoning about multiple interacting constraints, explaining the difficulty gap.
@@ H.4 Recommendations for Practitioners
    Apply difficulty-adaptive resources: Allocate more compute to Hard types (A, D, E, F, G; 59% avg RR@5) than Easy types (B, C; 90.5% avg).